# Unsupervised Constitutive Model Discovery from Sparse and Noisy Data

### v1

@@ 3.2 Constitutive Model Library
where $I_{1},I_{2},I_{3}$ are the standard invariants of $\boldsymbol{\mathbf{C}}$, see [38]. The isochoric component is modeled as a polynomial expansion:
@@ 3.3 Polynomial Chaos Expansion for Forecasted Displacements
In this work, uncertainty in forecasted displacements is introduced exclusively through the applied boundary tractions. The traction field $\overline{\boldsymbol{\mathbf{t}}}(\xi)$ from (13) is modeled as a random field with stochastic dependence on a scalar Gaussian random variable $\xi\sim\mathcal{N}(0,1)$. Specifically, we define:
@@
where $\bm{\mu}_{\overline{\boldsymbol{\mathbf{t}}}}$ denotes the deterministic mean traction vector and $\bm{\sigma}_{\overline{\boldsymbol{\mathbf{t}}}}$ is the standard deviation vector. To propagate this uncertainty through the forward problem, there exist intrusive [39, 40] and non-intrusive approaches [41, 42]. We employ a non-intrusive Polynomial Chaos (PC) expansion to approximate the stochastic displacement field. The forecasted (prior) displacement $\mathbf{u}^{\text{f}}(\xi)$ is then expressed as:
@@
where $\mathbf{u}_{j}\in\mathbb{R}^{n_{\text{gdof}}}$ are deterministic PC coefficients with $n_{\text{gdof}}=n_{\text{dof}}\cdot d$ the total number of degrees of freedom and $d$ the number of dimensions. Moreover, $\Psi_{j}(\xi)$ are Hermite polynomials, and $P_{u}$ is the total number of modes in the expansion.
In this contribution, the stochastic regression method is chosen to identify the PC coefficients $\mathbf{u}_{j}$ [43]. Given $S$ Monte Carlo samples $\{\xi_{s}\}_{s=1}^{S}$ and the corresponding finite element solutions $\mathbf{u}(\xi_{s})$, the coefficients are computed by minimizing the mean squared error between the actual displacements and their PC surrogate:
@@
where $\langle(\Psi_{j})^{2}\rangle$ denotes the expected value of the squared basis function. Note that because the single random variable $\xi$, the covariance matrix $\boldsymbol{\mathbf{C}}_{\mathbf{u}^{\text{f}}}$ exhibits nonzero off-diagonal terms, which reflects spatial correlation of the displacement field. The PC approximation $\mathbf{u}^{\text{f}}(\xi)$ serves as the prior displacement field in the subsequent Bayesian update, where it is assimilated with noisy sensor observations through the statFEM, see Sec. 4.
@@ 4 Statistical Inference with statFEM
$\mathbf{y}=\boldsymbol{\mathbf{H}}\,\mathbf{u}^{\mathrm{f}}+\mathbf{e}$, (26)
with $\boldsymbol{\mathbf{H}}\in\mathbb{R}^{n_{\text{gsen}}\times n_{\text{gdof}}}$ with $n_{\text{gsen}}=n_{\text{sen}}\cdot d$ is the linear observation matrix projecting full-field displacements to sensor locations, c.f. [30, 33, 35] for details. We assume that the forecasted displacement $\mathbf{u}^{\text{f}}$ follows a Gaussian distribution obtained from the polynomial chaos expansion (Sec. 3.3):
@@
$\int_{\Omega}\boldsymbol{\mathbf{P}}(\nabla\mathbf{u};\bm{\kappa}):\nabla\mathbf{v}\,\mathrm{d}\Omega=\int_{\partial\Omega_{p}}\overline{\mathbf{t}}_{\!N}\cdot\mathbf{v}\,\mathrm{d}S+\int_{\Omega}\mathbf{b}\cdot\mathbf{v}\,\mathrm{d}\Omega$. (32)
@@
$\bm{\kappa}_{\lambda}=\arg\min_{\bm{\kappa}\geq\mathbf{0}}\left\|\boldsymbol{\mathbf{A}}(\mathbf{y})\cdot\bm{\kappa}-\mathbf{p}\right\|_{2}^{2}+\lambda\mathbf{1}^{\top}\bm{\kappa}\quad\text{subject to}\,\,\mathbf{u}-\mathbf{y}=\mathbf{0}$. (34)
Here, the regularization term is written as $\lambda\mathbf{1}^{\top}\bm{\kappa}$, which is equivalent to the standard LASSO penalty $\lambda\|\bm{\kappa}\|_{p=1}$ for $\bm{\kappa}\geq\mathbf{0}$. The non-negativity constraint $\bm{\kappa}\geq\mathbf{0}$ is imposed as a sufficient condition for physical admissibility, ensuring that the identified strain energy function remains convex in the chosen basis. We note, however, that this is not a strictly necessary condition—other sign patterns may still be physically consistent depending on the form of the basis functions and the specific material behavior under consideration [44].
@@
that is, $n^{\mathcal{A}}_{\lambda}$ counts how many coefficients of the solution vector $\bm{\kappa}_{\lambda}$ are greater than $10^{-10}$ at a specific $\lambda$ such that they can be assumed to equal zero numerically. Additionally, we define the admissible set of the penalty terms as
@@
Crucially, model selection is not solely about reducing error, but about representing with interpretability and generalization capability [45]. To this end, we first minimize the model complexity, measured by the number of active coefficients $n^{\mathcal{A}}_{\lambda}$, i.e.,
@@ 6 Numerical Examples
This section demonstrates the performance of the proposed statFEM-EUCLID framework through synthetic numerical experiments. We consider the plate with a hole depicted in Fig. 3 as a case study.
@@
Synthetic Data Generation. Two benchmark materials are used as ground truth: (i) a Neo–Hookean model with one isochoric and one volumetric parameter (Sec. 6.1), and (ii) a Mooney–Rivlin model with two isochoric and one volumetric parameter (Sec. 6.2). The ground-truth displacement fields $\mathbf{u}^{\text{true}}=\mathbf{u}(\bm{\kappa}^{\text{true}})$ are computed by solving the deterministic forward FE problems with the true strain energy density $W^{\text{true}}$ and a right-edge traction of the form
@@
statFEM-EUCLID Iterative Discovery. The proposed iterative framework is initialized with a naive linear elastic prior. In both cases, we consider
@@ Acknowledgement
Upon acceptance of this manuscript, the source code supporting this study will be made available on GitHub and Zenodo [46].
@@ Appendix B Maximum a posteriori estimation of state and material parameters
$\mathbf{r}_{\mathrm{e}}(\mathbf{u},\mathbf{y})$ $=\boldsymbol{\mathbf{O}}(\mathbf{u})-\mathbf{y}$, (62)
@@ Appendix C The Nearest Symmetric Positive Definite Matrix
Numerical realization (Higham’s method): A practical and robust way to evaluate $\operatorname{SPD}(\mathbf{M})$ is based on the symmetric polar decomposition of the symmetrized matrix [47]. The following steps compute the optimizer of (68) up to machine precision:

### v2

@@ 3.2 Constitutive Model Library
where $I_{1},I_{2},I_{3}$ are the standard invariants of $\boldsymbol{\mathbf{C}}$, see [kim2014introduction]. The isochoric component is modeled as a polynomial expansion:
@@ 3.3 Polynomial Chaos Expansion for Forecasted Displacements
In this work, uncertainty in forecasted displacements is introduced exclusively through the applied boundary tractions. This choice provides a simple stochastic prior for the statFEM update and keeps the formulation as lightweight as possible. The proposed framework does not rely on this specific assumption. Uncertainty may be introduced in the material parameters, in the external loading, or in both. In the present study we focus on the simplest stochastic setting in order to demonstrate the iterative reconstruction and discovery procedure. More general stochastic formulations can be incorporated in future work. The traction field $\overline{\boldsymbol{\mathbf{t}}}(\xi)$ from (13) is modeled as a random field with stochastic dependence on a scalar Gaussian random variable $\xi\sim\mathcal{N}(0,1)$. Specifically, we define:
@@
where $\bm{\mu}_{\overline{\boldsymbol{\mathbf{t}}}}$ denotes the deterministic mean traction vector and $\bm{\sigma}_{\overline{\boldsymbol{\mathbf{t}}}}$ is the standard deviation vector. To propagate this uncertainty through the forward problem, there exist intrusive [deb2001solution, matthies2005galerkin] and non-intrusive approaches [berveiller2006stochastic, reagana2003uncertainty]. We employ a non-intrusive PCE to approximate the stochastic displacement field. The forecasted (prior) displacement $\mathbf{u}^{\text{f}}(\xi)$ is then expressed as:
@@
where $\mathbf{u}_{j}\in\mathbb{R}^{n_{\text{gdof}}}$ are deterministic PC coefficients with $n_{\text{gdof}}=n_{\text{dof}}\cdot d$ the total number of degrees of freedom and $d$ the number of dimensions. Moreover, $\Psi_{j}(\xi)$ are Hermite polynomials, and $P_{u}$ is the total number of modes in the expansion. The use of a PCE surrogate reduces the computational cost of the Bayesian update to negligible levels during the iterative discovery process, as no finite element solves are required in the data-assimilation step.
In this contribution, the stochastic regression method is chosen to identify the PCE coefficients $\mathbf{u}_{j}$ [berveiller2005non]. Given $S$ Monte Carlo samples $\{\xi_{s}\}_{s=1}^{S}$ and the corresponding finite element solutions $\mathbf{u}(\xi_{s})$, the coefficients are computed by minimizing the mean squared error between the actual displacements and their PCE surrogate:
@@
where $\langle(\Psi_{j})^{2}\rangle$ denotes the expected value of the squared basis function. Note that because the single random variable $\xi$, the covariance matrix $\boldsymbol{\mathbf{C}}_{\mathbf{u}^{\text{f}}}$ exhibits nonzero off-diagonal terms, which reflects spatial correlation of the displacement field. The PCE approximation $\mathbf{u}^{\text{f}}(\xi)$ serves as the prior displacement field in the subsequent Bayesian update, where it is assimilated with noisy sensor observations through the statFEM, see Sec. 4.
@@ 4 Statistical Inference with statFEM
$\mathbf{y}=\boldsymbol{\mathbf{H}}\,\mathbf{u}^{\mathrm{f}}+\mathbf{e}$. (26)
We assume that the forecasted displacement $\mathbf{u}^{\text{f}}$ follows a Gaussian distribution obtained from the PCE (Sec. 3.3):
@@
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\int_{\Omega}\boldsymbol{\mathbf{P}}(\nabla\mathbf{u};\bm{\kappa}):\nabla\mathbf{v}\,\mathrm{d}\Omega-\int_{\partial\Omega_{\boldsymbol{\mathbf{P}}}}\overline{\mathbf{t}}_{\!N}\cdot\mathbf{v}\,\mathrm{d}S=\mathbf{0}.}$ (32)
@@
$\bm{\kappa}_{\lambda}=\arg\min_{\bm{\kappa}\geq\mathbf{0}}\left\|\boldsymbol{\mathbf{A}}(\mathbf{y})\bm{\kappa}-\mathbf{p}\right\|_{2}^{2}+\lambda\mathbf{1}^{\top}\bm{\kappa}\quad\text{subject to}\,\,\boldsymbol{\mathbf{H}}\,\mathbf{u}-\mathbf{y}=\mathbf{0}$. (34)
Here, the regularization term is written as $\lambda\mathbf{1}^{\top}\bm{\kappa}$, which is equivalent to the standard LASSO penalty $\lambda\|\bm{\kappa}\|_{p=1}$ for $\bm{\kappa}\geq\mathbf{0}$. The non-negativity constraint $\bm{\kappa}\geq\mathbf{0}$ is imposed as a sufficient condition for physical admissibility. Since the basis functions in the library are chosen to be polyconvex with respect to the deformation measures, enforcing non-negative coefficients ensures that the identified total strain energy density preserves these stability properties [hartmann_generalizedStrainEnergyFunctions_2003]. We note, however, that this is not a strictly necessary condition—other sign patterns may still be physically consistent depending on the form of the basis functions and the specific material behavior under consideration [ricker_systematicFittingHyperelasticity_2023].
@@
that is, $n^{\mathcal{A}}_{\lambda}$ counts how many coefficients of the solution vector $\bm{\kappa}_{\lambda}$ are greater than $10^{-10}$ at a specific $\lambda$ such that the rest can be assumed to equal zero numerically. Additionally, we define the admissible set of the penalty terms as
@@
Crucially, model selection is not solely about reducing error, but about representing with interpretability and generalization capability [brunton2022data]. To this end, we first minimize the model complexity, measured by the number of active coefficients $n^{\mathcal{A}}_{\lambda}$, i.e.,
@@ 6 Numerical Examples
This section demonstrates the performance of the proposed statFEM-EUCLID framework through synthetic numerical experiments. We acknowledge that experimental validation is the ultimate goal; however, in this paper, synthetic data is utilized to allow for an objective quantification of the discovery error against a known ground truth, which is not possible with experimental data where the true constitutive law is unknown. Synthetic data allows us to systematically vary noise levels and sensor configurations to verify the algorithm’s convergence properties, which is the primary focus of this methodological contribution.
@@
Synthetic Data Generation. We consider the plate with a hole depicted in Fig. 3 as a case study. Two benchmark materials are used as ground truth: (i) a Neo–Hookean model with one isochoric and one volumetric parameter (Sec. 6.1), and (ii) a Mooney–Rivlin model with two isochoric and one volumetric parameter (Sec. 6.2). The ground-truth displacement fields $\mathbf{u}^{\text{true}}=\mathbf{u}(\bm{\kappa}^{\text{true}})$ are computed by solving the deterministic forward FE problems with the true strain energy density $W^{\text{true}}$ and a right-edge traction of the form
@@
statFEM-EUCLID Iterative Discovery. The proposed iterative framework is initialized with a naive linear elastic material model. It is used to compute forecasted displacements in the first iteration of the framework. While a more sophisticated constitutive model could be employed if expert knowledge were available, we deliberately select this simple, model-agnostic baseline to demonstrate the robustness of the discovery algorithm. In both test cases, we consider
@@ Acknowledgement
Upon acceptance of this manuscript, the source code supporting this study will be made available on GitHub and Zenodo [knauf_narouie_2025_17249520].
@@ Appendix B Maximum a posteriori estimation of state and material parameters
[…]
@@ Appendix C The Nearest Symmetric Positive Definite Matrix
Numerical realization (Higham’s method): A practical and robust way to evaluate $\operatorname{SPD}(\mathbf{M})$ is based on the symmetric polar decomposition of the symmetrized matrix [higham1988computing]. The following steps compute the optimizer of (68) up to machine precision: